OpenStudy (anonymous):

So, In Calc 1 we are reviewing some elementry function type stuff, would anyone like to help me?

OpenStudy (anonymous):

post!

OpenStudy (anonymous):

please post

OpenStudy (anonymous):

What is the difference between a natural domain and a regular domain?

OpenStudy (anonymous):

am i asking this in the wrong spot?

OpenStudy (anonymous):

no just post a question. like find \[\cos(\sin^{-1}(x))\] or is \[f(x)=\frac{x}{x^2-1}\] even odd or neither

OpenStudy (anonymous):

okay, so using the question you just posted as an example.. how would i figure out whether its even or odd or neither

OpenStudy (zarkon):

evaluate f(-x)

OpenStudy (anonymous):

Sorry to sound like an idiot but to evaluate simplify the denominator?

OpenStudy (anonymous):

in the function replace all the x with -x if it comes out to be -x then it is an odd function

OpenStudy (anonymous):

oh, okay. thanks.

OpenStudy (anonymous):

i must be psychic

OpenStudy (amistre64):

a natural domain is one that is defined for the function; such as sqrt(x) has a natural domain that is defined to make it work